# Accurate Models of NVIDIA Tensor Cores

### v1

@@ Abstract
The software associated with this paper, the MATLAB Tensor Core v0.1, which includes various NVIDIA GPU tensor core models as well as a generalised model that can be used to instantiate custom tensor core variants, is available on GitHub: https://github.com/north-numerical-computing/MATLAB-tensor-core.
@@ IV-A1 V100 Data Center GPU
Mixed-precision matrix multiplication units, or tensor cores, were introduced in the NVIDIA Volta architecture, namely, the V100 device. The V100 tensor cores support only fp16 as the input format with fp16 or fp32 as the output format [27]. Xi et al. [11] claim that the FMA size cannot be determined when there are no extra alignment bits compared with the precision of the output format. Following [13] we denote this situation with ${n_{\mathrm{eab}}}=0$. However, using the FMA size detection algorithm proposed in [13], we were able to identify an FMA size of $4$, which is valid for both cases of ${n_{\mathrm{eab}}}=0$ and ${n_{\mathrm{eab}}}\neq 0$. In addition, when the significands are aligned, bits that fall outside the internal accumulator’s precision are truncated rather than accumulated into multiple sticky bits as is done in the algorithm by Tenca [25]. Further tests on the V100 GPU using this approach led us to the estimated tensor core model shown in Fig. 2, which differs from the model presented in [28, Fig. 2] in that the significand alignment is performed in 25 bits rather than 24, and $c_{1}$ is added together with the products rather than separately.
@@ IV-A2 A100 Data Center GPU
The approximate model is shown in Fig. 3 (a). For tf19 input mode, the FMA size is reduced to $4$, with truncation as the default rounding mode. The bit-width of the denormalised sum must be $29$-bit wide. The model is shown in Fig. 3 (b). The products remain denormalised for fp16/bf16/tf19 input modes, as previously discussed in the context of V100. This behaviour can be verified by setting $c=0$, $p_{1}=2.25$ where $a_{1}=b_{1}=1.5$, and $p_{2}=2^{-23}$, $p_{3}=p_{4}=2^{-24}$. When $d=2.25+2^{-22}$, the products are added in denormalised form. The difference in this test compared to the V100 lies in the presence of an extra alignment bit (${n_{\mathrm{eab}}}=1$) in the A100, which slightly alters the accumulation behaviour, making this test dependent on knowing the $n_{\mathrm{eab}}$} value.
@@
In fp64 input/output mode, the FMA size is $1$. This mode is compliant with IEEE 754 with RNE as the default rounding mode, but supports RD, RU and RZ. The observed accumulation order appears to follow
@@
IV-A3 H100 Data Center GPU
Furthermore, the FMA size is $16$, which equals the maximum supported shared dimension of the matrix sizes available in the WMMA API. By the same reasoning, the denormalized adder result has a width of $27+\left\lceil\log_{2}(17)\right\rceil=32$ bits. Upon normalization, results are truncated rather than rounded. The fp16 format for both the input and output is supported, and the default rounding mode from the internal precision to the fp16 output is RNE. For the tf19 input mode, all numerical features are identical to those of the A100 model, except that two extra alignment bits are used instead of one. The fp64 input/output mode exhibits numerical features consistent with the fp64 tensor core of the A100.
The H100 and H200 GPUs also support two fp8 formats as the input format in the tensor cores, but the WMMA API does not provide direct support for those formats. Therefore, MMA instructions, which are PTX-level instructions, must be invoked directly. The matrix sizes supported by the MMA instruction for the fp8 input formats are
$m\times n\times k=16\times 8\times 16\quad\text{and}\quad 16\times 8\times 32.$ (2)
Here $k$ denotes the shared dimension of the input $A$ and $B$ matrices supported by the tensor cores and can either be $k=16$ or $32$. For both fp8 floating-point formats, fp8-E5M2 and fp8-E4M3, the output can be obtained either in fp16 or fp32 format.
  • fp32 output: In this format, the inner product of two 32-element fp8 vectors is computed using two block FMAs of size 16. The first partial inner product $s_{1}$ is computed as
    $s_{1}=\sum_{\ell}p_{\ell},\quad\ell=1,2,5,6,9,10,\dots,29,30.$
    During this partial computation, two extra alignment bits beyond the 23 fractional bits are used, while the rest of the lower order bits are truncated. The accumulation then proceeds in single precision, with $s_{1}$ subsequently normalized and truncated to fp32. In this process, denormalized products during accumulation may occur, similar to the behavior observed for fp16, bf16, and tf19 input formats. For the second half of the inner product of 32 elements, the normalized and truncated $s_{1}$ is used as an addend in computing
    $s_{2}=s_{1}+\sum_{\ell}p_{\ell},\quad\ell=3,4,7,8,11,12,\dots,31,32.$
    Here, the same features are present, except that alignment may be relative to $s_{1}$ if it is the largest operand, unlike in the first case where the input was zero and only the products were accumulated. Finally, the normalized and truncated result (in fp32) is added to $c$ using the default round-to-nearest-even (RNE) rounding mode.
    This interleaved pattern for computing the inner product of 32-element vectors can be verified by fixing $p_{1}=1$ and $p_{2}=2^{-24}$, and then assigning the value $2^{-24}$ sequentially to $p_{3}$ through $p_{32}$ while keeping all other entries at zero. If $d=1+2^{-23}$, the tested product falls into the same accumulation group as $p_{1}$ and $p_{2}$. Conversely, if $d=1$, the tested product belongs to the second group. If the two groups were not interleaved, we would observe $d=1+2^{-23}$ only when assigning $2^{-24}$ to positions $p_{3}$$p_{16}$, and $d=1$ when assigning it to positions $p_{17}$$p_{32}$. For $k=16$ case, we detected the same interleaved pattern.
  • fp16 output: For fp16 outputs $c$ and $d$, the internal accumulation still takes place in fp32 plus the extra alignment and carry bits. The first partial sum $s_{1}$ is normalized and rounded to fp16 using RNE. It is then used to compute $s_{2}$ in fp32 accumulation which is again normalized and reduced to fp16 precision with RNE. Finally, $c$ is added to the reduced and normalized $s_{2}$, with the final rounding mode set to RNE.
The FMA size detection algorithms presented previously [11, 13] are not directly applicable here, as they are constrained to the WMMA API. Furthermore, applying algorithms therein or test vectors directly, without modification, results in incorrect FMA size detection due to the overlapped nature of the halving of the 32-element vector into two parts. The best estimate of the tensor core model for the H100 GPU is illustrated in Fig. 4. It is also important to note that the fp8 model characteristics of the H100 GPU differ significantly from those of the Ada Lovelace RTX 1000 and NVIDIA L40S GPUs as discussed below.
Figure 4: A model of the inner product within the tensor cores of the H100/H200/B200 GPUs for (a) fp16/BF16, (b) tf19, and (c) fp8 input formats. The fp64 tensor core is compliant with IEEE 754 FMA operation and is not shown.
The NVIDIA H200 GPU tensor cores exhibit numerical characteristics identical to those of the H100 (Fig. 4).
IV-A4 B200 Data Center GPU Model
The B200 tensor cores, upon testing, reveal identical numerical behaviour to the H100/H200 models (Fig. 4). For single-precision accumulation with fp16/bf16 or tf19 inputs, the accumulator provides $2$ extra bits, and the FMA tile sizes are $16/16$ and $4$, respectively. The fp8 tensor core also manifests the same interleaved inner-product partitioning of $32$-element vectors as discussed previously for the H100/H200.
IV-A5 L40S Data Center GPU Model
Running the GNFT for fp16, bf16, and tf19 input formats, we determined that the L40S tensor cores exhibit behaviour identical to that of the A100, with FMA sizes of $8$, 88, and $4$, respectively, and with a single extra alignment bit when aligning significands, giving a 24-bit alignment/accumulation fraction in total. Bits that fall off the 24-bit fraction in the alignment of the significands during the multi-term floating-point addition are truncated, and the final results are truncated as well, unless the output is required in fp16 for the fp16 input setting. In this case, RNE is used as the default rounding mode.
When the input format is set to one of the fp8 formats, the interleaved pattern present for $k=32$ in the H100/B200/H200 is not present in the L40S or RTX-1000 Ada (a consumer-grade GPU that we had available on a laptop and have also tested). For both fp8 formats, fp8-E5M2 and fp8-E4M3, the accumulation takes place with $13$ fractional bits. We determined this by setting $p_{1}=1$, $p_{2}=p_{3}=2^{-23-{n_{\mathrm{eab}}}}$ and decrementing $n_{\mathrm{eab}}$} from $10$ until $d$ become equal to $1+p_{1}+p_{2}$. For ${n_{\mathrm{eab}}}=-10$, we obtain $d=1+2^{-12}$, which indicates that there are 2310=1323-10=13 fractional bits, unlike the H100/H200/B200 where accumulation takes place within the 25-bit fraction. Moreover, we determined that $c$ is added early with the first $16$ products by setting $c=1$ and $p_{1}=p_{2}=2^{-14}$, which produced $d=1$ in fp32 output mode.
The Ada Lovelace RTX 1000 consumer-grade GPU tensor cores exhibited identical numerical characteristics to those of the L40S model. This is further supported by the identical outputs produced by both models’ tensor cores for inputs drawn from a single ensemble of $10^{5}$ randomized input vectors. This is most likely because both GPUs belong to the Ada family and have fourth generation tensor cores on board. Since only the fp8 tensor core structure of the L40S is different compared with the tensor cores in the Ampere architecture (which does not accommodate fp8 at all), its best estimate is shown in Figure 5 only for the fp8.
Figure 5: A model of the inner product within the tensor cores of the L40S and Ada Lovelace RXT 1000 GPU for the fp8 input format.
IV-A6 A2 & A30 Data Center GPUs
@@
Input Output GPU architectures Prd. Align. Bitwidth Acc. out. prec. $N_{\mathrm{FMA}}$} Final rounding $c$ ord. in acc. Interleaved inputs
fp8 fp32 H100, H200, B200 (2,25) 32 16 RNE late yes
@@
Note: subnormal in/out supported; fp16 output for fp16 input is supported on all mentioned GPUs (rounded via RNE). The products remain denormalised in alignment and accummulation, reflected via $2$ integer bits. Extra alignment bits $n_{\mathrm{eab}}$} are included in the fractional bits of the product alignment precision. Accumulation output precision is a sum of extra carry bits and the product alignment bits.
@@ IV-B Authentication Via Randomized Testing
For the hardware execution, the WMMA API was used for all supported data types except fp8, for which MMA instructions were invoked in CUDA. To further verify that tensor cores were active, the cuobjdump tool was used to inspect the compiled binary and confirm the presence of HMMA/QMMA/DMMA instructions. Across all tested configurations, the MATLAB models produced results that matched the hardware outputs exactly at the bit level for every run. It is important to note that this exact agreement was only achieved after properly incorporating the handling of denormalized products into the software models. A mismatch within the loop of Algorithm 1 was instrumental in revealing this feature and reiterating the validation. Furthermore, the interleaved pattern used in computing the inner products of input vectors on the H100/H200/B200 tensor cores was identified through a combination of discrepancies between software and hardware outputs for identical inputs and the results of the $N_{\mathrm{FMA}}$} determination algorithm in [13]. This algorithm produced ${N_{\mathrm{FMA}}}=2$, indicating an interleaved pattern in the inner product computation; otherwise, $N_{\mathrm{FMA}}$} would not be smaller than ${N_{\mathrm{FMA}}}=16$ for fp16/bf16. In short, the refinement Algorithm 1 was executed for at least one iteration to find new features of tensor cores.
@@ IV-C Example MATLAB Code for Using the Models
A direct example of CustomTC model is provided in the form of B200TCRN which uses B200 features but with params.frmode set to rne. There is also a field stkbitenabled in params, i.e., params.stkbitenabled, which, when the significands of product terms are aligned, appends an extra sticky bit beyond the extra alignment bits, i.e., $n_{\mathrm{eab}}$}. By default, this is set to 0. In addition, the field params.inter_pattern is included due to the H100/H200/B200 tensor core behavior with the fp8 input format, discussed in Section IV-A3. If this field is set to $1$, the custom model will compute the inner product of $2{N_{\mathrm{FMA}}}$-element vectors in two interleaved vectors, and the variable $c$ is then added to the inner product at the end using a fixed RNE rounding mode (see Figure 4). The default value of this field is 0.
The GEMM.m file executes the parallelized version of matrix multiplication if the Parallel Computing Toolbox, introduced in Matlab R2013b, is installed and the machine supports multicore processing. This ensures that computations are efficiently distributed across available CPU cores, accelerating large-scale matrix operations. If either is not supported, serialized version is executed.

### v2

@@ Abstract
The software associated with this paper, the MATLAB Tensor Core v0.2, which includes various NVIDIA GPU tensor core models as well as a generalised model that can be used to instantiate custom tensor core variants, is available on GitHub: https://github.com/north-numerical-computing/MATLAB-tensor-core.
@@ IV-A1 V100 Data Center GPU
Mixed-precision matrix multiplication units, or tensor cores, were introduced in the NVIDIA Volta architecture, namely, the V100 device. The V100 tensor cores support only fp16 as the input format with fp16 or fp32 as the output format [27]. Xi et al. [11] claim that the FMA size cannot be determined when there are no extra alignment bits compared with the precision of the output format. Following [13] we denote this situation with ${n_{\mathrm{eab}}}=0$. However, using the FMA size, $N_{\mathrm{FMA}}$}, detection algorithm proposed in [13], we were able to identify ${N_{\mathrm{FMA}}}=4$, which is valid for both cases of ${n_{\mathrm{eab}}}=0$ and ${n_{\mathrm{eab}}}\neq 0$. In addition, when the significands are aligned, bits that fall outside the internal accumulator’s precision are truncated rather than accumulated into multiple sticky bits as is done in the algorithm by Tenca [25]. Further tests on the V100 GPU using this approach led us to the estimated tensor core model shown in Fig. 2, which differs from the model presented in [28, Fig. 2] in that the significand alignment is performed in 25 bits rather than 24, and $c_{1}$ is added together with the products rather than separately.
@@ IV-A2 A100 Data Center GPU
The approximate model is shown in Fig. 3 (a). For tf19 input mode, the $N_{\mathrm{FMA}}$} is $4$, with truncation as the default rounding mode. The bit-width of the denormalised sum must be $29$-bit wide. The model is shown in Fig. 3 (b). The products remain denormalised for fp16/bf16/tf19 input modes, as previously discussed in the context of V100. This behaviour can be verified by setting $c=0$, $p_{1}=2.25$ where $a_{1}=b_{1}=1.5$, and $p_{2}=2^{-23}$, $p_{3}=p_{4}=2^{-24}$. When $d=2.25+2^{-22}$, the products are added in denormalised form. The difference in this test compared to the V100 lies in the presence of an extra alignment bit (${n_{\mathrm{eab}}}=1$) in the A100, which slightly alters the accumulation behaviour, making this test dependent on knowing the $n_{\mathrm{eab}}$} value.
@@
In fp64 input/output mode, we have ${N_{\mathrm{FMA}}}=1$. This mode is compliant with IEEE 754 with RNE as the default rounding mode, but supports RD, RU and RZ. The observed accumulation order appears to follow
@@
IV-A3 A2 & A30 Data Center GPUs
@@
IV-A4 L40S Data Center GPU Model
Running the GNFT for fp16, bf16, and tf19 input formats, we determined that the L40S tensor cores exhibit behaviour identical to that of the A100, with $N_{\mathrm{FMA}}$} of $8$, 88, and $4$, respectively, and with a single extra alignment bit when aligning significands, giving a 24-bit alignment/accumulation fraction in total. Bits that fall off the 24-bit fraction in the alignment of the significands during the multi-term floating-point addition are truncated, and the final results are truncated as well, unless the output is required in fp16 for the fp16 input setting. In this case, RNE is used as the default rounding mode.
The WMMA API does not provide support for fp8 format tensor cores available in Ada architecture. Therefore, the MMA instruction, specifically mma.sync.aligned.m16n8k16/32.f32/f16.f8.f8.f32/f16, is utilized which then maps to QMMA instruction, for computing fp8 MMA across a warp [29, Sec. 6.4] which we found is also present in Ada architecture, albeit not mentioned in its instruction set [29, Sec. 6.2]. For both fp8 formats, fp8-E5M2 and fp8-E4M3, the accumulation takes place with $13$ fractional bits. We determined this by setting $p_{1}=1$, $p_{2}=p_{3}=2^{-23-{n_{\mathrm{eab}}}}$ and decrementing $n_{\mathrm{eab}}$} from $10$ until $d$ becomes equal to $1+p_{1}+p_{2}$. For ${n_{\mathrm{eab}}}=-10$, we obtain $d=1+2^{-12}$, which indicates that there are 2310=1323-10=13 fractional bits. Moreover, we determined that $c$ is added early with the first $16$ products by setting $c=1$ and $p_{1}=p_{2}=2^{-14}$, which produced $d=1$ in fp32 output mode.
The Ada Lovelace RTX 1000 consumer-grade GPU tensor cores exhibited identical numerical characteristics to those of the L40S model. This is further supported by the identical outputs produced by both models’ tensor cores for inputs drawn from a single ensemble of $10^{5}$ randomized input vectors. This is most likely because both GPUs belong to the Ada family and have fourth generation tensor cores on board. Since only the fp8 tensor core structure of the L40S is different compared with the tensor cores in the Ampere architecture (which does not accommodate fp8 at all), its best estimate is shown in Figure 4.
Figure 4: A model of the inner product within the tensor cores of the L40S and Ada Lovelace RXT 1000 GPU for the fp8 input format. For fp16, bf16 and tf19, the model is identical to A100 and is not shown.
IV-A5 H100 Data Center GPU
Furthermore, the $N_{\mathrm{FMA}}$} for fp16 and bf16 is $16$, which equals the maximum supported shared dimension of the matrix sizes available in the WMMA API. By the same reasoning, the denormalized adder result has a width of $27+\left\lceil\log_{2}(17)\right\rceil=32$ bits. Upon normalization, results are truncated rather than rounded. The fp16 format for both the input and output is supported, and the default rounding mode from the internal precision to the fp16 output is RNE.
Using the WMMA API with the only available shape for the input matrix, i.e., m16n16k8, for tf19 input format, the instructions are mapped to HMMA but with fragment size of m16n8k4 which results in an FMA size i.e., $N_{\mathrm{FMA}}$}, of $4$ via GNFT. However, if mma instruction is used with matrix shape of m16n8k8, it is internally mapped to HMMA however with the size m16n8k8 which shows that the $N_{\mathrm{FMA}}$} of $8$ if the GNFT is applied. We report the maximum $N_{\mathrm{FMA}}$} in the estimate model of H100 and H200 (see Figure 5).
With WMMA API limited to fp16, bf16, tf19 and binary64 formats, mma instruction, as for Ada RTX 1000 and L40S, can be used to multiply fp8 matrices. However, unlike Ada RTX 1000 or L40S where this is mapped to QMMA, in the case of H100 and H200, initially a conversion takes place from fp8 to fp16, and then HMMA is utilised, with the matrix shape of m16n8k16. This shows that direct QMMA instruction is not available in Hopper architecture and fp8 inputs with mma are simply multiplied with fp16 tensor cores. Moreover, the fp8 inner product, with the mma is computed in interleaved fashion, with $c$ added via RNE (see Figure. 5c, where 32-element input vector is distributed across two tensor core invocations by alternating pairs of elements). The interleaved input pattern is detected in both cases for m16n8k16 and m16n8k32 matrix shapes. It can be verified by fixing $p_{1}=1$ and $p_{2}=2^{-24}$, and then assigning the value $2^{-24}$ sequentially to $p_{3}$ through $p_{32}$ while keeping all other entries at zero. If $d=1+2^{-23}$, the tested product falls into the same accumulation group as $p_{1}$ and $p_{2}$. Conversely, if $d=1$, the tested product belongs to the second group. If the two groups were not interleaved, we would observe $d=1+2^{-23}$ only when assigning $2^{-24}$ to positions $p_{3}$$p_{16}$, and $d=1$ when assigning it to positions $p_{17}$$p_{32}$.
The Hopper architecture supports a specific warpgroup level mma instruction i.e., wgmma.mma, which is internally mapped to QGMMA, responsible for fp8 MMA across a warpgroup (a set of four contiguous warps) [29, Sec. 6.3], instruction at assembly level which supports native fp8 format tensor core access. With this, the GNFT shows ${n_{\mathrm{eab}}}=-10$, i.e., $13$ fractional bits and an FMA size of $32$. The structure of fp8 format tensor core for H100/H200 is identical to that of the Ada Lovelace architecture, but with twice the $N_{\mathrm{FMA}}$}. The model diagram for H100 and H200 is shown in Figure 5.
Figure 5: A model of the inner product within the tensor cores of the H100/H200/B200 GPUs for (a) fp16/BF16, (b) tf19, (c) fp8 input format provided via mma.sync (which internally uses fp16 tensor core with interleaved input pattern) and (d) fp8 tensor core accessed via wgmma.mma_async (specific to Hopper architecture). The fp64 tensor core is compliant with IEEE 754 FMA operation and is not shown.
IV-A6 B200 Data Center GPU Model
The B200 tensor cores, upon testing, reveal identical numerical behaviour to the H100/H200 models (Fig. 5). For single-precision accumulation with fp16/bf16 or tf19 inputs, the accumulator provides $2$ extra bits, and the FMA tile sizes are $16/16$ and $8$, respectively. We test the fp8 tensor core only through the mma PTX instruction and we found identical behaviour to the H100/H200, where fp8 input vectors are converted to fp16 and HMMA instruction is called. The interleaving of pairs of values from the input vectors was also present for both $k=16$ and k=32k=32—the model is shown in Figure 5(c). Even though QMMA is specifically mentioned in the instruction set of Blackwell GPU [29, Sec. 6.4], our experiments show that mma.sync instructions are mapped to HMMA not QMMA. This confirms that QMMA requires PTX instructions other than mma. This observation of mma.sync not mapping to QMMA but instead to HMMA is consistent with the findings of [30].
@@ IV-A6 B200 Data Center GPU Model
Input Output GPU architectures Prd. Align. Bitwidth Acc. out. prec. $N_{\mathrm{FMA}}$} Final rounding $c$ ord. in acc. Interleaving
fp8 fp32 H100, H200 (2,13) 21 32 Trunc early no
@@
Note: subnormal in/out supported; fp16 output for fp16 input is supported on all mentioned GPUs (rounded via RNE). The products remain denormalised in alignment and accummulation, reflected via $2$ integer bits. Extra alignment bits $n_{\mathrm{eab}}$} are included in the fractional bits of the product alignment precision. Accumulation output precision is a sum of extra carry bits and the product alignment bits. * fp8 implementation with mma.sync.aligned.m16n8k16/32.f32/f16.f8.f8.f32/f16 where fp8 to fp16 conversion takes place and HMMA instructions are called.
@@ IV-B Authentication Via Randomized Testing
For the hardware execution, the WMMA API was used for all supported data types except fp8, for which MMA instructions were invoked in CUDA. To further verify that tensor cores were active, the cuobjdump tool was used to inspect the compiled binary and confirm the presence of HMMA/QMMA/QGMMA/DMMA instructions. Across all tested configurations, the MATLAB models produced results that matched the hardware outputs exactly at the bit level for every run. It is important to note that this exact agreement was only achieved after properly incorporating the handling of denormalized products into the software models. A mismatch within the loop of Algorithm 1 was instrumental in revealing this feature and reiterating the validation. Furthermore, the interleaved pattern used in computing the inner products of input vectors on the H100/H200/B200 tensor cores was identified through a combination of discrepancies between software and hardware outputs for identical inputs and the results of the $N_{\mathrm{FMA}}$} determination algorithm in [13]. This algorithm produced ${N_{\mathrm{FMA}}}=2$, indicating an interleaved pattern in the inner product computation; otherwise, $N_{\mathrm{FMA}}$} would not be smaller than ${N_{\mathrm{FMA}}}=16$ for fp16/bf16. In short, the refinement Algorithm 1 was executed for at least one iteration to find new features of tensor cores.
@@ IV-C Example MATLAB Code for Using the Models
A direct example of CustomTC model is provided in the form of B200TCRN which uses B200 features but with params.frmode set to rne. There is also a field stkbitenabled in params, i.e., params.stkbitenabled, which, when the significands of product terms are aligned, appends an extra sticky bit beyond the extra alignment bits, i.e., $n_{\mathrm{eab}}$}. By default, this is set to 0. In addition, the field params.inter_pattern is included due to the H100/H200/B200 tensor core behavior with the fp8 input format, discussed in Section IV-A5. If this field is set to $1$, the custom model will compute the inner product of $2{N_{\mathrm{FMA}}}$-element vectors in two interleaved vectors, and the variable $c$ is then added to the inner product at the end using a fixed RNE rounding mode (see Figure 5). The default value of this field is 0.
The GEMM.m file executes the parallelized version of matrix multiplication if the Parallel Computing Toolbox, introduced in Matlab R2013b, is installed and the machine supports multicore processing. This ensures that computations are efficiently distributed across available CPU cores, accelerating large-scale matrix operations. If either is not supported, serialized version is executed. Lastly, the proposed toolbox is compatible with Octave and can also be accessed from Python using either Oct2Py library or the Matlab engine.